OpenStudy (anonymous):

Rectangle ABCD is similar to Rectangle EFGH. If AB = 5 feet, BC = 16 feet, and FG = 21 feet, what is the length of EF? 6 9/16 correct?

OpenStudy (binary3i):

AB/EF = BC/FG 5/x = 16/21 x=105/16

OpenStudy (anonymous):

6 + 9/16 = 105 / 16. So you were right.

OpenStudy (anonymous):

oqei thanks